OpenStudy (anonymous):

Find the discriminant and determine the real natural roots 2x^2-4x+2=0

jhonyy9 (jhonyy9):

2(x-1)(x-1)=2(x^2 -2x +1) =2x^2 -4x +2 x_1 =x_2 = 1

OpenStudy (anonymous):

Discriminant is b^2-4ac for ax^2 +bx +c so -4^2 - 4*2*2 = 0 Zero discriminant means repeated roots (of 1, in this case).